Suppose your credit card balance is ​$17000 . The minimum payment is ​$347​, and the annual percentage rate is 20.1​%.

a.  If you make a constant monthly payment of ​$347, how long will it take you to pay off the credit card​ balance?

b.  How much interest will you pay if you elect to make the minimum​ payment?

Please complete using Excel and show all formulas used.
Answer #1

Answer :

a. To pay off Credit card balance with monthly payment $347 , it takes 104 months i.e 8.6years.

b. Total interest to be paid if we elect to pay minimum amount is $18892.23.

Calculations in excel format is shown below :

Answer a.

Interest = Balance amount *1.675% (where interest rate /12 = 20.1%/12 = 1.675%) i.e 1.675% is monthly interest rate

Principle = Payment amount - Interest amount

Balance = Previous month balance - Principle amount

Calculation for 1st month

Interest =17000*1.675% = 284.75

Principle = 347-284.75 = 62.25

Balance = 17000 - 62.25 = 16937.75

follow the same calculation for every month.

Month Payment Interset Principle Balance
17000
1 347 284.75 62.25 16937.75
2 347 283.7073 63.29269 16874.46
3 347 282.6472 64.35284 16810.1
4 347 281.5692 65.43075 16744.67
5 347 280.4733 66.52672 16678.15
6 347 279.359 67.64104 16610.51
7 347 278.226 68.77403 16541.73
8 347 277.074 69.92599 16471.81
9 347 275.9027 71.09725 16400.71
10 347 274.7119 72.28813 16328.42
11 347 273.501 73.49896 16254.92
12 347 272.2699 74.73006 16180.19
13 347 271.0182 75.98179 16104.21
14 347 269.7455 77.25449 16026.96
15 347 268.4515 78.5485 15948.41
16 347 267.1358 79.86419 15868.54
17 347 265.7981 81.20191 15787.34
18 347 264.438 82.56204 15704.78
19 347 263.055 83.94496 15620.83
20 347 261.649 85.35104 15535.48
21 347 260.2193 86.78067 15448.7
22 347 258.7658 88.23424 15360.47
23 347 257.2878 89.71217 15270.76
24 347 255.7852 91.21484 15179.54
25 347 254.2573 92.74269 15086.8
26 347 252.7039 94.29613 14992.5
27 347 251.1244 95.87559 14896.63
28 347 249.5185 97.48151 14799.14
29 347 247.8857 99.11432 14700.03
30 347 246.2255 100.7745 14599.26
31 347 244.5375 102.4625 14496.79
32 347 242.8213 104.1787 14392.61
33 347 241.0763 105.9237 14286.69
34 347 239.3021 107.6979 14178.99
35 347 237.4981 109.5019 14069.49
36 347 235.664 111.336 13958.16
37 347 233.7991 113.2009 13844.95
38 347 231.903 115.097 13729.86
39 347 229.9751 117.0249 13612.83
40 347 228.0149 118.9851 13493.85
41 347 226.0219 120.9781 13372.87
42 347 223.9956 123.0044 13249.86
43 347 221.9352 125.0648 13124.8
44 347 219.8404 127.1596 12997.64
45 347 217.7105 129.2895 12868.35
46 347 215.5449 131.4551 12736.9
47 347 213.343 133.657 12603.24
48 347 211.1043 135.8957 12467.34
49 347 208.828 138.172 12329.17
50 347 206.5136 140.4864 12188.68
51 347 204.1605 142.8395 12045.85
52 347 201.7679 145.2321 11900.61
53 347 199.3353 147.6647 11752.95
54 347 196.8619 150.1381 11602.81
55 347 194.3471 152.6529 11450.16
56 347 191.7901 155.2099 11294.95
57 347 189.1904 157.8096 11137.14
58 347 186.5471 160.4529 10976.69
59 347 183.8595 163.1405 10813.54
60 347 181.1269 165.8731 10647.67
61 347 178.3485 168.6515 10479.02
62 347 175.5236 171.4764 10307.54
63 347 172.6514 174.3486 10133.19
64 347 169.731 177.269 9955.926
65 347 166.7618 180.2382 9775.688
66 347 163.7428 183.2572 9592.43
67 347 160.6732 186.3268 9406.104
68 347 157.5522 189.4478 9216.656
69 347 154.379 192.621 9024.035
70 347 151.1526 195.8474 8828.187
71 347 147.8721 199.1279 8629.059
72 347 144.5367 202.4633 8426.596
73 347 141.1455 205.8545 8220.742
74 347 137.6974 209.3026 8011.439
75 347 134.1916 212.8084 7798.631
76 347 130.6271 216.3729 7582.258
77 347 127.0028 219.9972 7362.261
78 347 123.3179 223.6821 7138.578
79 347 119.5712 227.4288 6911.15
80 347 115.7618 231.2382 6679.911
81 347 111.8885 235.1115 6444.8
82 347 107.9504 239.0496 6205.75
83 347 103.9463 243.0537 5962.697
84 347 99.87517 247.1248 5715.572
85 347 95.73583 251.2642 5464.308
86 347 91.52715 255.4728 5208.835
87 347 87.24798 259.752 4949.083
88 347 82.89714 264.1029 4684.98
89 347 78.47341 268.5266 4416.453
90 347 73.97559 273.0244 4143.429
91 347 69.40243 277.5976 3865.831
92 347 64.75268 282.2473 3583.584
93 347 60.02503 286.975 3296.609
94 347 55.2182 291.7818 3004.827
95 347 50.33086 296.6691 2708.158
96 347 45.36165 301.6384 2406.52
97 347 40.30921 306.6908 2099.829
98 347 35.17214 311.8279 1788.001
99 347 29.94902 317.051 1470.95
100 347 24.63842 322.3616 1148.589
101 347 19.23886 327.7611 820.8274
102 347 13.74886 333.2511 487.5763
103 347 8.166903 338.8331 148.7432
104 347 2.491448 344.5086 -195.765
Total int. 18892.23

Answer b.

Add all amounts in interest column by using SUM formula i.e =Sum(number 1,number 2,..,Number n) , which is $18892.23.